What is the distance fallen after 4 s for a freely falling object starting from rest ? Use g = 9.8 m/s2. State your final answer in a numerical form, up to the first decimal place, with appropriate units in an abbreviated form.
Answer #1

Use second equation of motion

S = ut +1/2 at^2

Starting from rest u=0

S= 0 +1/2x 9.8 x 4x4 = 78.4m